An order for a patient reads “Give 40. mg of pantopazole IV and 5 g of MgSO IV.” The pantopazole should be administered at a concentration of 0.4 mg/mL and the MgSO should be administered at a concentration of 0.02 g/mL in separate IV infusion bags. What is the total fluid volume the patient has received from both IV infusions?

clear handwriting and step by step explanation please and please
Answer #1

Mass of pantopazole IV given = 40 mg

Mass of MgSO IV = 5 g

Concentration of pantopazole = 0.4 mg/mL

Concentration of MgSO = 0.02 g/mL

Volume of pantopazole given

= mass of pantopazole / Concentration of pantopazole

= (40 mg) / (0.4 mg/mL)

V1 = 100 mL

Volume of MgSO given

= mass of MgSO / Concentration of MgSO

= (5 g) / (0.02 g/mL)

V2 = 250 mL

total fluid volume the patient has received from both IV infusions = V1 + V2

= 100 + 250 = 350 mL